# An Efficient Algorithm for Permutation Iteration Using a Singly Linked List

### v3

@@ 2 Description of the algorithm
We refer to these recursive calls as “left” or “right”, depending on the position of the previously popped element. The general structure of the algorithm can now be illustrated by the following triangle:
@@ 3 Proof of correctness
To understand how the sublist is modified by a recursive call, let us examine the last permutation produced when iterating over all permutations of $(0,1,2,\dots,k-1)$:
@@
Lemma 3.1.
Let $k$ be an odd number $(k>4)$, and assume that the last permutation of $k$ elements produced by the algorithm is the initial sequence with the first two elements swapped. When the algorithm is applied to the sequence $(0,1,2,\dots,k)$, each of the $k+1$ elements will be inserted exactly once at the beginning, and the final permutation generated by the algorithm will be $(1,4,3,5,6,7,8,\dots,k,2,0)$.
@@
Lemma 3.2.
Let’s $k$ be an even number $(k>4)$, and assume that the last permutation of $(0,1,2,\dots,k-1)$ produced by the algorithm is $(1,4,3,5,6,7,8,\dots,k-1,2,0)$. When the algorithm is applied to the sequence $(0,1,2,\dots,k)$, each of the $k+1$ elements will be inserted exactly once at the beginning, and the final permutation generated will be the initial sequence with the first two elements swapped.
@@
Next, $k$ “left” recursive calls are made. However, tracking the position of each element throughout the process is not as straightforward as in the proof of Lemma 3.1. An simpler approach can be achieved by leveraging graph theory.
@@
The algorithm works correctly for $k$ in the range from 00 to $5$. Additionally, we verify that the last computed permutation of $(e_{0},e_{1},e_{2},e_{3},e_{4})$ is $(e_{1},e_{0},e_{2},e_{3},e_{4})$. Using Lemmas 3.1 and 3.2, we prove by induction that the algorithm selects each element of the sequence exactly once to insert it at the beginning before computing all permutations of the remaining elements. ∎
@@ 4 Transition Distance Between Consecutive Permutations
$\lim_{k\to\infty}\frac{D_{k}}{k!-1}=7\cosh 1-4\sinh 1-5\approx 1.100759669\dots$
@@ 5 Implementation technical details
We now discuss an implementation of the algorithm using singly linked lists. Such a data structure is typically not well-suited for extracting elements repeatedly from both the beginning and the end of a sublist. However, in this very specific case, a natural and elegant solution to this issue emerges.

### v4

@@ 2 Description of the algorithm
We refer to these recursive calls as “left” or “right”, depending on the position of the previously popped element (more precisely, each of these terms denotes the two sequential operations: the list modification and the following recursive call). The general structure of the algorithm can now be illustrated by the following triangle:
@@ 3 Proof of correctness
To understand how the sublist is modified during a recursive call, consider the last permutation produced while iterating over all permutations of $(0,1,2,\dots,k-1)$. Although the primary purpose of the recursive call is to generate multiple permutations, it can also be viewed externally as a direct mapping of $(0,1,2,\dots,k-1)$ to a specific resulting permutation:
@@
In accordance with the pattern inherent in the algorithm, two distinct cases can be easily identified based on the parity of $k$, which will now be examined in two corresponding lemmas.
Lemma 1.
Let $k$ be an odd number $(k>4)$, and assume that the last permutation of $k$ elements produced by the algorithm turns out to be the initial sequence with the first two elements swapped. When the algorithm is applied to the sequence $(0,1,2,\dots,k)$, each of the $k+1$ elements will be inserted exactly once at the beginning, and the final permutation generated by the algorithm will be $(1,4,3,5,6,7,8,\dots,k,2,0)$.
@@
Lemma 2.
Let’s $k$ be an even number $(k>4)$, and assume that the last permutation of $(0,1,2,\dots,k-1)$ produced by the algorithm turns out to be $(1,4,3,5,6,7,8,\dots,k-1,2,0)$. When the algorithm is applied to the sequence $(0,1,2,\dots,k)$, each of the $k+1$ elements will be inserted exactly once at the beginning, and the final permutation generated will be the initial sequence with the first two elements swapped.
@@
Next, $k$ “left” recursive calls are made. However, tracking the position of each element throughout the process is not as straightforward as in the proof of Lemma 1. An simpler approach can be achieved by leveraging graph theory.
@@
The algorithm works correctly for $k$ in the range from 00 to $5$. Additionally, we verify that the last computed permutation of $(e_{0},e_{1},e_{2},e_{3},e_{4})$ is $(e_{1},e_{0},e_{2},e_{3},e_{4})$. Using Lemmas 1 and 2, we prove by induction that the algorithm selects each element of the sequence exactly once to insert it at the beginning before computing all permutations of the remaining elements. ∎
@@ 4 Transition Distance Between Consecutive Permutations
$\lim_{k\to\infty}\frac{D_{k}}{k!-1}=7\cosh(1)-4\sinh(1)-5\quad\approx 1.100759669\dots$
@@
As a final summary, each permutation is derived from the previous one by selecting an element from the current list and inserting it to the left, with an average movement of slightly more than one position.
@@ 5 Implementation technical details
We now discuss an implementation of the algorithm using singly linked lists. Such a data structure is typically not well-suited for extracting elements repeatedly from both the beginning and the end of a sublist. However, in this very specific case, a natural and elegant solution to this issue will emerge. We highlight that making this extraction an atomic operation is one of the key features of our algorithm.